OpenStudy (anonymous):

A water tanks has a diameter of 6m and depth of 5m. If the water is 3m deep and is rising at 5m/h, at what rate is the volume changing?

OpenStudy (anonymous):



geerky42 (geerky42):

Because we are only given rate of change in height, you would need to know the relationship between r and h, so you can make some substitutions so \(V\) would only contain one variable; \(h\) We are given that r = 3m when h = 5m. then we can find the relationahip using proportion: So we have \(\dfrac{r}{3} = \dfrac{h}{5}\Longrightarrow r = \dfrac{3}{5}h\) Now, it is known that \(V = \dfrac{1}{3}\pi r^2h\) Substitute \(r\) with \(\dfrac{3}{5}h\) and we have \(V = \dfrac{1}{3}\pi\left(\dfrac{3}{5}h\right)^2h = \dfrac{3}{25}\pi h^3\) Now we can take derivative of \(V\) with respect to \(t\) (time): \(\dfrac{dV}{dt} = \dfrac{9}{25}\pi h^2 \dfrac{dh}{dt}\) Now you are given \(h = 3~m\) and \(\dfrac{dh}{dt}=5~m/h\) (Positive because it is rising) So you can plug in these values and solve for \(\dfrac{dV}{dt}\) Does that help?